# {cutout}03.2cm55cm1 empty PDF-WuKong         : A Large Multimodal Model for Efficient Long PDF Reading with End-to-End Sparse Sampling

### v1

@@ 3.1 Overview
3.2 Document Parsing
3.3 Sparse Sampling
For the parsed $n$ text chunks $\{T_{1},T_{2},\ldots,T_{n}\}$, $m$ images $\{I_{1},I_{2},\ldots,I_{m}\}$, and the input user query $q$, we first extract the positive samples and the negative samples for the query. Our PaperPDF dataset has provided corresponding positive single-evidence or multi-evidence samples for each query-answer pair (detailed in Sec. 4). We randomly select two text blocks and two images from the remaining text blocks and images as negative samples. Then, we use a text encoder $En\_T$ to obtain the text embeddings $e_{T_{P}},e_{T_{N}}$ and the query embedding $e_{q}$. An image encoder $En\_I$ is utilized to output the image features $e_{I_{P}},e_{I_{N}}$, which is shared with MLLM.
$\mathcal{L}_{\text{rep}}=-\frac{1}{P}\sum_{e_{i}\in E_{P}}\log\frac{e^{{\text{sim}(e_{q},e_{i})}/\tau}}{e^{{\text{sim}(e_{q},e_{i})}/\tau}+\displaystyle\sum_{e_{j}\in E_{N}}e^{{\text{sim}(e_{q},e_{j})}/\tau}},$ (1)
where $\text{sim}(e_{q},e_{i})$ and $\text{sim}(e_{q},e_{j})$ represent the similarity between the query and the positive/negative samples, respectively. $\tau$ is the temperature parameter that controls the scale of the similarity scores. $P$ is the number of positive samples. By maximizing this probability, the model encourages the representations of the query and positive samples to be closer while pushing the representations of the query and negative samples apart. We align the feature space of the text encoder to the pre-trained vision encoder with frozen parameters. It is worth noting that this sparse sampler is a plug-and-play design that can be applied to any MLLMs.
During the inference, we pre-encode all text blocks and images and cache all candidate vector embeddings. When the user inputs a query, we calculate the similarity between query embedding and cached text/image embeddings. Then the model automatically selects the top-$k$ relevant text blocks and images as evidence to respond to this query. Therefore, this process samples out sparse document content, greatly reducing the computational burden of the subsequent LLM and alleviating the problem of attention shift when facing ultra-long sequences. Moreover, the multimodal embedding cache further optimizes inference time.
3.4 Answer Generation
At this stage, the large language model only receives the document content that is most relevant to the query and discards a lot of redundant information, so it can generate more accurate answers with higher efficiency. Specifically, we input the sampled top-$k$ evidence, the user query, and the task instruction into the LLM, and let it generate an answer based on the provided query and evidence. The inference process is shown in Algorithm 1.
PDF-WuKong is optimized end-to-end by these two loss functions for effective multimodal alignment and question-answering. The training pipeline of PDF-WuKong is shown in Algorithm 2.
4 PaperPDF Construction
4.1 Overview
The reasons for constructing our dataset include the following two points. In most long document Q&A scenarios, the basis for an answer to a given question is typically composed of a single element or a combination of multiple elements, while the remaining information may act as noise, interfering with the reasoning of MLLMs. Additionally, existing document question-answering datasets are either limited to single-page formats or provide only page-level ground truth, which hinders the training of the sparse sampler in PDF-WuKong. Consequently, we propose a reliable method for generating high-quality question-answer pairs of long document and develop the dataset PaperPDF for training and evaluation based on this method.
Category Training Test
Single Text-only 249k 2939
Image-only 21K 212
Multi Image-text 250k 2566
Section 499k 255
Cross-paragraph 1.2k 118
Table 2: The statistic of PaperPDF. Text-only and Image-only indicate that the QA pairs are generated based on either a single text paragraph or an image extracted from the PDF. Image-text, Section, and Cross-paragraph denote that the QA pairs are generated from a paragraph and its corresponding references, an entire section, or non-contiguous paragraphs, respectively.
4.2 Detail of Construction
The construction process of PaperPDF are presented in Fig. 3. It can be divided into four steps: structured parsing, rule-based extraction, prompt construction, and filtering. We obtained 89k PDF documents from the arXiv repository as our document set $\mathbb{D}$. For each document $D$ in $\mathbb{D}$, we first employed Grobid to parse the document and extract its text chunks $\{T_{1},T_{2},\ldots,T_{m}\}$ and images chunks $\{I_{1},I_{2},\ldots,I_{n}\}$. Subsequently, predefined rules are used to randomly select chunks from the document, which may consist of text chunks $T_{i}$, chart chunks $I_{j}$ or a combination of both. The selected chunks will be input into commercial MLLM products according to different prompt templates and then, the question $Q$ related to the input chunks, along with the corresponding answer $A$, is subsequently generated. Notably, For the training set, we used Gemini for generation due to its free and rapid accessibility, while the high-performance GPT-4V was employed to construct the test set, ensuring the validity and robustness of the evaluation. Finally, we devise a set of rules to filter the generated training and testing sets automatically. The removing rules for the samples include too-short questions, too-long answers, non-English text, etc. Further manual checking is conducted on test set to ensure the reliability of the evaluation. PaperPDF consists of two types of QA pairs: Single-evidence and Multi-evidence.
Single-evidence QA pair means that the question can be answered based on a single text chunk or image chunk in the long document. It can be systematically categorized into two types: text-only and image-only. In text-only generation, a text chunk is input into MLLM, whereas in image-only generation, an image chunk is utilized. The generation process for Single-evidence QA pairs is relatively straightforward and cost-effective, primarily aimed at training the foundational multimodal understanding and question-answering capabilities of MLLMs.
Multi-evidence QA pairs. Unlike single-evidence QA pairs, Multi-evidence QA data exhibits greater complexity, both in terms of characteristics and generation process. The answers to such questions typically rely on multiple text chunks, image chunks, or any combination thereof in document. Multi-evidence QA pairs consist of Image-text QA pairs, Section QA pairs, and Cross-paragraph QA pairs. The Image-text QA pairs are generated based on a paragraph and its corresponding image references. These pairs require consideration of the interrelationship between the text and the image to ensure the answers capture multimodal information. Section QA pairs are designed to train the MLLM in integrating and understanding information within a section and are generated from all chunks within a section. Cross-paragraph QA pairs involve the most complex generation process. First, the entire document is input into the MLLM for paragraph-level semantic summarization, followed by the selection of semantically related text chunks. Finally, multiple related chunks are randomly selected and re-input into the MLLM to generate the final QA pairs. This type of data primarily focuses on training the sampler’s performance, enhancing its sparse sampling capability. Although the generation of Multi-evidence QA pairs is relatively complex and may require multiple MLLM calls, their presence significantly improves the performance of the sampler and strengthens the MLLM’s ability to comprehend cross-chunk information.
Totally, we generated 1.5M samples for training and 8K samples for testing. After filtering, the final dataset consists of 1M training data points and 6K test data points for subsequent training and evaluation. The statistics of PaperPDF are presented in Tab. 2.
Figure 4: Examples from PDF-WuKong and other proprietary products
5 Experiments
[…]
Model #param ANLS F1 ROUGE Token Latency GPT Acc
[…]
Closed Qwen [46] - 36.0 40.3 35.5 - - 78.1
Kimi [47] - 28.5 33.6 31.1 - - 74.7
Gemini pro [48] - 26.6 29.0 29.8 - - 67.9
ChatGLM [49] - 31.2 35.4 32.0 - - 73.5
[…]
5.1 Implementation Details
[…]
Table 4: Description of the input format used in the experiment
Symbol Description
* Input the parsed content of the pdf file.
[…]
$\dagger$ Input the entire page image.
$\ddagger$ Input the entire page image and OCR content.
5.2 Long PDF Understanding
To assess the effectiveness of our model in understanding long PDF documents, we conducted comprehensive experiments comparing it with both open-source models and commercial products on PaperPDF dataset.
Due to the limited availability of open-source models capable of handling this task, we firsly evaluate two such models that can process multi-page PDF documents. Considering that IXC2-VL [45] cannot understand such long image-text interleaved sequences, we also report the results with the input of pure OCR content. To further demonstrate the advantages of our multi-modal sampler, we demonstrate the evaluation of IXC2-VL with RAG where we feed the top 5 text paragraphs retrived by BGE-M3 [50]. The results on the PaperPDF dataset are reported in Tab. 3. The findings indicate that our model significantly outperforms existing open-source models across various evaluation metrics. Additionally, thanks to the introduction of the sparse sampler, the number of tokens that our model’s MLLM needs to process is substantially fewer than that of other models.
Moreover, we compared our model with several commercial products that allow users to input PDF documents and questions via web interfaces to generate answers. Considering the cost of manual testing, we randomly selected 50 PDFs from the PaperPDF dataset for evaluation, and the results are presented in Tab. 3. Our model clearly outperforms mainstream PDF question-answering products. Fig. 4 shows some examples from PDF-WuKong and other proprietary products. Our model can accurately retrieve the evidence needed to answer user questions and generate accurate answers based on relevant text segments or diagrams. Since these commercial solutions do not provide detailed code or technical reports, we could not obtain information about their parameter sizes or token counts, and uploading documents directly to their websites does not allow us to accurately assess the models’ latency. Finally, we evaluated the GPT accuracy of all open-source and proprietary models on this subset benchmark. The results show that our model has significant advantages in both textual similarity and semantic understanding.
To compare with more open-source large document models, considering that most of these models can only handle single-page documents, we constructed a subset of the PaperPDF benchmark containing only test samples with single evidence. Therefore, we provided all models with only the page containing the evidence as input; note that the pages are input in the form of images. These models and our PDF-WuKong have not been trained on such data. In this setting, our sparse sampler sampled all the content of that page. As shown in Tab.5, our model’s zero-shot capability on this subset is significantly better than other document large models.
Table 5: Performance comparison with other DocVLMs for PDF multimodal understanding on Single-Evidence Subset of PaperPDF. ($\dagger$) indicates the use of a special input approach. For details, refer to Table 4.
Model # param ANLS F1 ROUGE
[…]
5.3 Document-oriented VQA
To validate the generalization capability of our model to other document understanding scenarios, we conducted experiments on several benchmark datasets and compared PDF-WuKong with other representative models.
First, we evaluated the performance of PDF-WuKong and other open-source models on DocVQA [35], ChartQA [37], and InfoVQA [40], which are all single-page document datasets. As shown in Tab. 6, our model achieved leading performance compared to other open-source models. This demonstrates that PDF-WuKong can effectively handle various types of documents and questions, showcasing its versatility in document-oriented visual question answering tasks.
In addition, we assessed the performance of traditional specialized models and large-scale models on two existing multi-page document QA datasets. The experimental results, presented in Tab. 7, indicate that our model’s performance in multi-page document scenarios is comparable to these specialized models and far surpasses the latest document large model, DocOwl2 [26]. Notably, on complex multi-page document datasets like DUDE [58], PDF-WuKong outperforms GPT-4V [59]. This improvement is attributed to our sparse sampler, which effectively filters out useful information from multi-page documents, enabling the model to focus on relevant content.
Furthermore, we conducted zero-shot evaluations on a new long multimodal document understanding benchmark MM-NIAH [18]. As shown in Tab. 8, our model uses the fewest parameters yet achieves the second-best performance. Although InternVL-V1-5-RAG [18] surpasses PDF-WuKong by 2.8%, it utilizes 36.5 billion more parameters than our model. Moreover, as the context length of the multimodal documents increases, the performance of our model remains stable, unlike other models that experience significant declines. At a context length of 64K, PDF-WuKong achieves the best results, demonstrating its robustness in handling long-context multimodal inputs.
Table 6: Performance comparison with other DocVLMs on single-page document-oriented VQA benchmarks. ($\dagger$) indicates the use of a special input approach. For details, refer to Table 4.
Model Doc. Chart. Info.
Closed Gemini Pro [60] 88.1 74.1 75.2
GPT-4V [61] 88.4 78.5 75.1
[…]
Table 7: Performance comparison with other DocVLMs for multi-page document understanding. ($\dagger$) indicates the use of a special input approach. For details, refer to Table 4.
Model MP-DocVQA DUDE
[…]
Table 8: Performance comparison with other DocVLMs on MM-NIAH. The input approach aligns with the benchmark.
Model #param Overall 1K 4K 16K 64K
Emu2-Chat [55] 37B 8.8 38.9 18.2 0.0 0.0
VILA1.0-13b [69] 13B 15.7 41.9 33.2 8.6 0.1
llava-v1.6-13b [70] 13B 16.9 43.7 34.9 13.6 0.0
llava-v1.6-34b [23] 34B 20.6 57.4 45.1 8.2 0.0
InternVL1.5 [11] 26B 41.1 59.5 50.1 41.9 16.6
InternVL1.5-RAG [18] 45B 46.1 59.5 50.1 44.9 39.3
PDF-WuKong (ours) 8.5B 43.3 53.0 43.9 43.0 42.1
5.4 Ablation Study
To comprehensively evaluate the effectiveness of our proposed model components, we conducted ablation studies focusing on datasets, the impact of the sparse sampler, sampling strategies, and document length. Below, we present the findings from each of these experiments.
Datasets
We conducted experiments to evaluate the impact of training data size and composition on model performance. As shown in Tab. 9, increasing the amount of training data led to consistent improvements in the model’s accuracy on the PaperPDF benchmark. Specifically, the model trained with larger datasets demonstrated better comprehension and answer accuracy. Moreover, under the same data volume, training sets that included multi-evidence annotations outperformed those containing only single-evidence. This suggests that exposure to diverse and complex evidence during training enhances the model’s ability to handle queries requiring multiple pieces of information, thereby improving overall performance.
Table 9: Ablation study on dataset setting
Dataset ANLS F1 ROUGE
100 k 38.7 40.1 37.5
500 k 41.6 43.5 40.8
1 M 42.6 43.6 40.2
Impact of sparse sampler
To assess the effectiveness of the sparse sampler, we compared models trained with and without it. Without the sparse sampler, the MLLM struggled to process long documents with interleaved text and images, resulting in poor performance due to the overwhelming amount of irrelevant information. Introducing the sparse sampler significantly improved the model’s accuracy, as evidenced in Tab. 10, by efficiently selecting the most relevant content for each query. Furthermore, end-to-end joint training of the sparse sampler and the MLLM led to additional performance gains compared to training them separately. This indicates that our integrated design allows the two components to mutually enhance each other, optimizing both the retrieval of pertinent information and the generation of accurate answers.
Table 10: Ablation study on the impact of sparse sampler
Sparse Sampler End-to-End ANLS F1 ROUGE
11.1 5.1 5.0
40.3 42.3 39.8
42.6 43.6 40.2
Sampling strategy
We explored different sampling strategies by varying the number of top N text paragraphs or figures selected by the sparse sampler. As shown in Tab. 11, setting N too low resulted in missing critical information necessary for accurately answering queries, thereby reducing performance. Conversely, a larger N introduced redundant information and increased computational and time costs without significant improvements in accuracy. This suggests that there is an optimal range for N that balances the inclusion of essential information with efficiency. Selecting the appropriate number of samples is crucial for maximizing performance while minimizing resource consumption.
Table 11: Ablation study of different sampling strategy
Sampling chunks ANLS F1 ROUGE Tokens
Top 1 39.20 38.28 35.26 1186
Top 3 42.09 43.06 39.69 1452
Top 5 42.59 43.63 40.19 1789
Top 10 43.01 44.22 40.67 2386
Top 15 43.19 44.57 42.08 2704
Top 20 43.42 45.02 42.30 3364
Figure 5: Ablation study of different document length
Document length
To understand the impact of document length on model performance and efficiency, we divided the test set into subsets based on the number of pages per document. Results in Fig. 5 demonstrate that our model’s performance and time efficiency remained relatively stable across documents of varying lengths. This stability indicates that the sparse sampler effectively reduces the input size to a manageable level, regardless of the original document length. In contrast, MLLMs without the sparse sampler were unable to handle long documents effectively; their performance and time efficiency deteriorated significantly as the document length increased. These findings highlight the robustness of our approach in processing long documents without sacrificing accuracy or incurring additional computational costs.
6 Conclusion
We have presented PDF-WuKong, a novel Multimodal Large Language Model that effectively addresses the challenges of understanding long PDF documents containing interleaved text and images. By introducing an end-to-end sparse sampling mechanism, our model efficiently extracts the most relevant paragraphs and diagrams in response to user queries, significantly reducing input token size and making the process independent of document length. We also constructed PaperPDF, a comprehensive dataset with 1 million question-answer pairs for training and 6,000 pairs for evaluation, specifically tailored for academic PDFs. Experimental results demonstrate that PDF-WuKong not only outperforms existing open-source models but also surpasses proprietary products by an average of 8.6% in F1 score on long multimodal PDF understanding tasks. Our approach maintains high accuracy and efficiency even as document length increases, offering a scalable and interpretable solution for practical applications in document understanding.
Appendix A Dataset Examples
Text-only Data Example in the Training Set Query: What is the impact of using the same dataset for optimizing and measuring the performance of a model? Text: Here, again, the unfair advantage of optimizing (selecting the models for the ensemble) and measuring performance on the same dataset appears.The advantage is small but systematic for the test split of ISIC (Fig. 5a); it is much more apparent for the challenging collection of clinical images of EDRA Atlas (Fig. 5b). Answer 1: It can lead to an unfair advantage for the model. Answer 2: Optimizing a model involves selecting certain parameters or features that improve its performance on a given dataset. If the same dataset is used to measure the model’s performance, it may lead to an unfair advantage as the model has already been “tuned” to that specific dataset.
Text-image Data Example in the Training Set Query: What is the best performing method for both detection and classification, according to the provided figure? Text: 3) The last one is our proposed SFCN-OPI with both sibling branches and OPI (Ours in Table 1). Figure: Table 1[Uncaptioned image] Caption: Experimental results of ablation analysis, different training strategies of our method and comparison with other approaches. Note: the -means the results were not reported by that method. Answer 1: Ours. Answer 2: The best performing method for both detection and classification is Ours, as it achieves the highest F1 scores for both tasks. This can be seen in the “Ours” row of the table, where the F1 score for detection is 0.834 and the F1 score for classification is 0.742.
Section Data Example in the Training Set Query: What kind of information does the factor R i, j hold? Text:
RECOMMENDATION MODEL
In this section, we present our Joint Topic-Semantic-aware Matrix Factorization (JTS-MF) model for online social votings, in which social relationship, group a liation, and topic-semantic similarities are combined and taken into account for voting recommendation in a comprehensive manner.Motivated by Locally Linear Embedding [23] which tries to preserve the local linear dependency among inputs in the low-dimensional embedding space, we expect to keep inter-user and inter-voting topic-semantic similarities in latent feature space as well.To this end, in JTS-MF model, while the rating R i, j is factorized as user latent feature Q i and voting latent feature P j , we deliberately enforce Q i and P j to be dependent on their social-topic-semantic similar counterparts, respectively.e graphic model of JTS-MF model is as shown in Figure 5. Figure: Figure 5[Uncaptioned image] Caption: Fig. 5: Graphic Model of JTS-MF. Answer 1: The factor R i, j represents the rating that user i gave to voting j. Answer 2: The node labeled R i, j is a factor that represents the rating given by user i to a voting j. It is the core of the model, representing the essence of user interaction with the voting system. This node’s existence suggests that the model is primarily concerned with understanding and predicting user ratings for voting content.
Text-only Data Example in the Test Set Query: How does the distribution of output image transform tasks among N workers impact the efficiency of the final synchronization task? Text: The number of output image transform tasks equals the number of output images times the batch size. The tasks are executed by all N workers, such that each worker picks up an arbitrary task and executes it. The last output image transform task to finish also executes the final synchronization task, which frees the memory required for the output image transforms. Answer 1: The distribution allows for parallel processing, potentially speeding up the final synchronization task. Answer 2: Ensures faster task completion, synchronization efficiency relies on all tasks’ completion.
Text-image Data Example in the Test Set Query: What distinguishing features can be observed between ‘hit’ and ‘miss’ samples for real fingerprints in Fig. 5 across different benchmarks? Text: Images found in these benchmarks can be observed in Fig. 5 of Section V.As we can see, variability exists not only across modalities, but also within modalities.Moreover, it is rather unclear what features might discriminate real from spoofed images, which suggests that the use of a methodology able to use data to its maximum advantage might be a promising idea to tackle such set of problems in a principled way. Figure: Figure 5[Uncaptioned image] Caption: Fig. 5. Examples of hit and missed testing samples lying closest to the real-fake decision boundary of each benchmark. A magnified visual inspection on these images may suggest some properties of the problem to which the learned representations are sensitive. Answer 1: Hit’ samples for real fingerprints show clearer ridge patterns and quality compared to ‘miss’ samples throughout the benchmarks. Answer 2: Hit real fingerprints are clearer, more distinct; Miss real samples are blurred, less defined.
Only-image Data Example in the Test Set Query: Based on Figure 4, which fuzzing technique consistently leads in coverage across all benchmarks over the 24-hour period? Figure: Figure 4[Uncaptioned image] Caption: Figure 4: Number of benchmarks on which each technique has the lead in coverage at each hour. A benchmark is counted for multiple techniques if two techniques are tied for the lead. Answer 1: FairFuzz consistently leads in coverage across all benchmarks over the 24-hour period in Figure 4. Answer 2: FairFuzz, highest coverage benchmark count over 24 hours.
Section Data Example in the Test Set Query: How does the qualitative evaluation of extractive summarizers using word clouds elucidate the differences in content focus between the original documents and the summaries? Text:
Here we use word cloud representations to give an intuitive interpretation of the content in the generated extractive summarizers.We create word clouds for the two best methods in Section 5.3.In this paper, we used an online tool called WordItOut5 to generate the word cloud representations.In all word clouds presented in this paper, a filter is used to display only the words with minimum frequency of 2.
Figure 3 shows a word cloud made by the aggregation of all the summaries generated by the PKUSUMSUM-Centroid method.This gives a sense of the content in those summaries.For contrast, we also generate a word cloud for the original news articles without the content of the generated summaries.Specifically, common words are first removed completely and then the word clouds are built with frequencies of surviving words.In essence, this shows what information remains apart from the generated summaries. The images clearly show a contrast of content.The summary wordcloud shows “earthquake” as its most prominent word.The image of the articles show less focus.If viewed alone, the reader would not quickly infer the gist of the original content.Similarly, Figure 4 represents “Lead” method.And here we see an even more stark difference.6. Comparing Twitter and News Media information. Figure: Figure 3[Uncaptioned image] Caption: Figure 3: The word clouds representing summaries generated by PKUSUMSUM-Centroid method (left) and original documents without the content of those summaries (right). Answer 1: Word clouds highlight the prominent themes in summaries versus original texts by displaying relative word frequencies visually. Answer 2: Visual contrast in word frequency highlights content focus differences.
Cross-paragraph Data Example in the Test Set Query: How can we leverage the proposed EDO approach to optimize the selection of datasets for specific algorithms, thereby enhancing the overall performance and validity of the algorithms? Text:

### v2

@@ Appendix A Algorithm
Appendix B Visualization Results
B.1 Qualitative Comparison with Other Products
The qualitative comparison between PDF-WuKong and other proprietary products on the long PDF understanding task is presented in Fig. 5. PDF-WuKong achieves significant performance advantages over others such as Qwen [66], ChatGLM [67], Kimi [68], and Gemini-pro [69]. The success stems from its ability to effectively sample key evidence relevant to the question and its strong multimodal understanding capabilities. This observation further validates the effectiveness of our method in enhancing the long PDF understanding capabilities of MLLMs.
@@ B.1 Qualitative Comparison with Other Products
Figure 5: Qualitative comparison between PDF-WuKong with other proprietary products. The red box indicates the evidence that the correct answer depends on.
B.2 Visualization Results on Chinese Documents
In addition to the outstanding performance on English documents, PDF-WuKong also demonstrates remarkable capabilities on Chinese documents. Fig. 6 shows its qualitative performance on the Chinese long PDF document understanding task. It effectively samples relevant evidence within the PDF and demonstrates the ability to derive accurate answers based on the sampled evidence.
Appendix C More Dataset Details
The Q-E-A triplets in the PaperPDF dataset are derived from approximately 60k English and 10k Chinese documents, encompassing nearly 70 disciplines such as computer science, engineering, and materials science.
To comprehensively illustrate the characteristics of the PaperPDF dataset, we conduct a detailed statistical analysis. The key statistics of the source documents in PaperPDF are presented in Tab. 12. Compared to previous multi-page document datasets, PaperPDF not only includes a significantly larger number of documents but also features more pages and OCR tokens. More importantly, PaperPDF encompasses both Chinese and English documents with interleaved text and images. Our dataset provides an important innovation driver and a comprehensive benchmark for the development of the multimodal long PDF document understanding task.
For the five types of data (Text-only, Image-only, Image-text, Section, Cross-paragraph) in single evidence and multiple evidence, Fig. 7 - Fig. 11 show their prompt engineering and a corresponding example, respectively.
Figure 6: Examples of PDF-WuKong on Chinese documents. The red box indicates the evidence that the correct answer depends on.
MPdataset Document Language Page Image Token
DUDE [17] 5k En 6 - 1831
MP-Docvqa [46] 6k En 8 - 2026
PaperPDF (En) 60k En 25 12 11371
PaperPDF (Zh) 10k Zh 11 7 3413
PaperPDF 70k En+Zh 23 11 10234
Table 12: Statistics of the source documents, where Page, Image, and Token represent the average number of pages per document, the average number of images in the document, and the average number of OCR tokens per document, respectively.
Text-only Question Generation Prompt Task Definition: Create 2 academic questions from a given research paper paragraph. Requirements: Analyze the paragraph thoroughly,understanding its content including the study’s objectives, ethods, results,and conclusions. Focus on the paragraph,not the entire paper. If the paragraph lacks valid information,return ‘quit’. You should use English. Develop 2 questions that: Are no more than 30 words. Incorporate knowledge from the paragraph. Are answered by text instead of one of the multiple choices. Are elicit detailed responses supported by the text. Expected Output: (Return ‘quit’ directly if the paragraph lacks valid information.) [Q1]: question1 here [Q2]: question2 here
Text-only Answer Generation Prompt Task Definition: Answer a question based on the material given. Requirements: The answers should: Be comprehensive and cover all relevant aspects. Accurately reflect the paragraph’s information and insights. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, no more than 20 words. [thinking procedure]: [A2]: answer2 here, no more than one sentence.
Text-only Data Example Query: What is the impact of using the same dataset for optimizing and measuring the performance of a model? Text: Here, again, the unfair advantage of optimizing (selecting the models for the ensemble) and measuring performance on the same dataset appears.The advantage is small but systematic for the test split of ISIC (Fig. 5a); it is much more apparent for the challenging collection of clinical images of EDRA Atlas (Fig. 5b). Answer 1: It can lead to an unfair advantage for the model. Answer 2: Optimizing a model involves selecting certain parameters or features that improve its performance on a given dataset. If the same dataset is used to measure the model’s performance, it may lead to an unfair advantage as the model has already been “tuned” to that specific dataset.
Figure 7: Text-only Q-E-A triplets generation prompt and data example.
Image-only Question Generation Prompt Task Definition: Formulate 2 academic questions based on a provided figure or table from a research paper. Requirements: The questions must directly reference and integrate information presented in the image and its caption, ensuring a cohesive understanding of the content depicted. You should use English. Develop 2 questions that are: No more than 30 words. Specific to the unique data or details visible in the figures/tables and are answerable only based on the material without inferring or speculating on details not explicitly explained by the figures/tables. Not mentioning the label of the figure/table directly or use words like ‘from the figure/table’. Expected Output: [Q1]:[Q2]:
Image-only Answer Generation Prompt Task Definition: Answer a question based on an image and its caption from a research paper. Requirements: The answers should: Always use English. Not infer or speculate on details not explicitly explained by the figures/tables. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: …  [A1]: answer1 here, no more than 20 words. [thinking procedure]: …  [A2]: answer2 here, no more than one sentence.
Image-only Data Example Query: Based on Figure 4, which fuzzing technique consistently leads in coverage across all benchmarks over the 24-hour period? Figure: Figure 4Refer to caption Caption: Figure 4: Number of benchmarks on which each technique has the lead in coverage at each hour. A benchmark is counted for multiple techniques if two techniques are tied for the lead. Answer 1: FairFuzz consistently leads in coverage across all benchmarks over the 24-hour period in Figure 4. Answer 2: FairFuzz, highest coverage benchmark count over 24 hours.
Figure 8: Image-only Q-E-A triplets generation prompt and data example.
Text-image Question Generation Prompt Task Definition: Formulate 2 academic questions based on a provided paragraph, a figure or table from a research paper. Requirements: The questions should not rely on the accompanying text but only the figure / table. However, you may use the provided text to understand the figure / table. You should use English. Develop 2 questions that are: No more than 30 words. Based on a provided figure of a research paper, without relying on accompanying text for the questions. However, you may use the provided text to understand the figure. Not mentioning the label of the figure/table directly or use words like ‘from the figure/table’. Expected Output: [Q1]:[Q2]:
Text-image Answer Generation Prompt Task Definition: Answer a question based on the material given. Requirements: The answers should be: Always using English. Comprehensive and cover all relevant aspects, as presented in the provided text and figure. Accurately reflect the information and insights offered by the research paper. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: …  [A1]: answer1 here, no more than 20 words. [thinking procedure]: …  [A2]: answer2 here, no more than one sentence.
Text-image Data Example Query: What is the best performing method for both detection and classification, according to the provided figure? Text: 3) The last one is our proposed SFCN-OPI with both sibling branches and OPI (Ours in Table 1). Figure: Table 1Refer to caption Caption: Experimental results of ablation analysis, … Answer 1: Ours. Answer 2: The best performing method for both detection and classification is Ours, as it achieves the highest F1 scores for both tasks. This can be seen in the “Ours” row of the table, where the F1 score for detection is 0.834 and the F1 score for classification is 0.742.
Figure 9: Text-image Q-E-A triplets generation prompt and data example.
Section Question Generation Prompt Task Definition: Formulate 2 academic questions based on a section from a research paper. Requirements: Carefully read and comprehend the entire provided section of the research paper to ensure a thorough understanding of its content, including key points, findings, methodologies, and conclusions. You should Always use English. Develop 2 questions that are: No more than 30 words. Requiring an integration of information from all paragraphs and figures/tables in the section. Not mentioning the label of the figure/table directly or use words like ‘from the figure/table’. Not based on common knowledge or assumptions not supported by the figures and tables. Expected Output: [Q1]:[Q2]:
Section Answer Generation Prompt Task Definition: Answer a question based on the material given. Requirements: The answers should be: Always using English. Not inferring or speculating on details not explicitly explained by the figures/tables. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: …  [A1]: answer1 here, no more than 20 words. [thinking procedure]: …  [A2]: answer2 here, no more than one sentence.
Section Data Example Query: How does the qualitative evaluation of extractive summarizers using word clouds elucidate the differences in content focus between the original documents and the summaries? Text:
Here we use word cloud representations to give an intuitive interpretation of the content in the generated extractive summarizers…
Figure 3 shows a word cloud made by the aggregation of all the summaries generated by the PKUSUMSUM-Centroid method… The images clearly show a contrast of content… Figure: Figure 3Refer to caption Caption: Figure 3: The word clouds representing summaries generated by PKUSUMSUM-Centroid method (left) and original documents without the content of those summaries (right). Answer 1: Visual contrast in word frequency highlights content focus differences. Answer 2: Word clouds highlight the prominent themes in summaries versus original texts by displaying relative word frequencies visually.
Figure 10: Section Q-E-A triplets generation prompt and data example.
Cross-paragraph Question Generation Prompt Task Definition: Based on the selected paragraph from a research paper that share a thematic or conceptual connection, formulate an insightful, open-ended question. This question should reflect the shared themes or concepts of your selections and relate to the broader context of the research paper. Requirements: Ascertain the underlying connection among the paragraphs and the figures/tables(if provided). Subsequently, craft an insightful, open-ended question that encapsulates the identified themes or connections, aiming to foster analytical thinking and in-depth discussion on the subject matter of the paper. Expected Output: [Q]: [Your generated question based on the shared themes or information]
Cross-paragraph Answer Generation Prompt Task Definition: Given some selected paragraphs from a research paper, and ensuring that these paragraphs share a certain level of association, you are to answer a question that is related to the content of these selected paragraphs. Requirements: Craft the 2 answers for the question that: Are directly derived from the provided figure, excluding information not found within the material. Are comprehensive and cover all relevant aspects, as presented in the provided figure. Expected Output: [A1]: [Insert a concise answer here, no more than 20 words.] [A2]: [Insert a detailed answer here,including a detailed “thought chain” or reasoning process.]
Cross-paragraph Data Example Query: How can we leverage the proposed EDO approach to optimize the selection of datasets for specific algorithms, thereby enhancing the overall performance and validity of the algorithms? Text: